40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1049     Entry time: Wed Oct 15 17:40:50 2008
Author: rana 
Type: Update 
Category: PSL 
Subject: PMC Offset adjusted 
I set the PMC servo input offset: closed the MOPA shutter, zeroed the mixer output with the offset slider,
relocked everything, and set the nominal to the new value of -6 V.
ELOG V3.1.3-